OpenStudy (anonymous):

use the definition of the derivative f'(x)=lim x->delta x of (f(x+delta x)-f(x))/delta x of f(x)=2/(x-5)

OpenStudy (watchmath):

sorry , why I post that on your question? ?

OpenStudy (watchmath):

ignore that! there was a but on this software

OpenStudy (watchmath):

but =bug

OpenStudy (anonymous):

alright

OpenStudy (watchmath):

Then I oblige to answer your question :). For \(f(x)=2/(x-5)\). The expression become \[\frac{1}{\Delta x}(2/(x+\Delta x -5)-2/(x-5))=\frac{-2\Delta x}{\Delta x(x-5)(x+\Delta x-5)}=\frac{-2}{(x-5)(x+\Delta x-5)}\] As \(\Delta x\to 0\) we have \(-2/(x-5)^2\).

OpenStudy (anonymous):

for some reason i cant see past the second equals sign... :/

OpenStudy (watchmath):

Try to refresh the browser. Press F5.

OpenStudy (anonymous):

nope still the same thing. and i have a mac. so i dont have f5...

OpenStudy (watchmath):

See the attachment.